In watching a transverse wave go by, an observer notes that 13 crests go by in a time of 3.0 s. If the distance between two successive crest is measured to be 0.75 m, (a) what is the speed of the wave? (b) What is its period?


1
Expert's answer
2021-01-31T15:11:36-0500

(a) Let's first find the period of the wave:


"T=\dfrac{3\ s}{13\ crests}=0.23\ s."

Then, we can find the frequency of the wave:


"f=\dfrac{1}{T}=\dfrac{1}{0.23\ s}=4.35\ Hz."

Finally, we can find the speed of the wave:


"v=f\lambda=4.35\ Hz\cdot0.75\ m=3.26\ \dfrac{m}{s}."

Answer:

"v=3.26\ \dfrac{m}{s}, T=0.23\ s."